Project Proposals for 6.895

October 15, 2003

• Improving Cilk

– Adaptively Parallel Processor Allocation for Cilk Jobs (Kunal Agrawal and Siddhartha Sen)

– Fast Serial-Append File I/O Mode Support for Cilk (Alexandru Caracas)

– A Space-Efficient Global Scheduler for Cilk (Jason Hickey and Tyeler Quentmeyer)

– Automatic Conversion of Non Series-Parallel DAGs to Series Parallel

DAGs (Sajindra Jayasena and Sharad Ganesh)

• Transactional Cilk – Language-Level Complex Transactions (C. Scott Ananian)

– An Evaluation of Nested Concurrent Transactions (Sean Lie)

– Atomic Transactions in Cilk (Jim Sukha)

– Transactions in Cilk (Xie Yong)

• Non-Determinacy Detection – Linear Time Detection of Determinacy Races (Jeremy Fineman)

– Parallel Nondeterminator (He Yuxiong)

– Parallel Nondeterminator (Wang Junqing)

• Using Cilk – Accelerating Multiprocessor Simulation (Kenneth C. Barr)

– Parallelizing METIS (Zardosht Kesheff)

– Parallelizing Sorting (Paul Youn)

– Implement FIR filter in parallel using Cilk (Pham Duc Minh)

• Cache-Oblivious Algorithms

– Cache-Oblivious Sorting for Burrows-Wheeler Transform (Advait D.Karande and Sriram Saroop)

– New Cache-Oblivious Algorithms (Zhang Jiahui and Neel Kamal)

– Cache-Oblivious, Lock-Free Algorithms (Seth Gilbert)

Improving Cilk

• Adaptively Parallel Processor Allocation for Cilk Jobs (Kunal Agrawal and Siddhartha Sen)

• Fast Serial-Append File I/O Mode Support for Cilk (Alexandru Caracas)

• A Space-Efficient Global Scheduler for Cilk (Jason Hickey and Tyeler Quentmeyer)

• Automatic Conversion of Non Series-Parallel DAGs to Series Parallel DAGs (Sajindra Jayasena and Sharad Ganesh)

Transactional Cilk

• Language-Level Complex Transactions (C. Scott Ananian)

• An Evaluation of Nested Concurrent Transactions (Sean Lie)

• Atomic Transactions in Cilk (Jim Sukha)

• Transactions in Cilk (Xie Yong)

Non-Determinacy Detection

• Linear Time Detection of Determinacy Races (Jeremy Fineman)

• Parallel Nondeterminator (He Yuxiong)

• Parallel Nondeterminator (Wang Junqing)

Using Cilk

• Accelerating Multiprocessor Simulation (Kenneth C. Barr)

• Parallelizing METIS (Zardosht Kesheff)

• Parallelizing Sorting (Paul Youn)

• Implement FIR filter in parallel using Cilk (Pham Duc Minh)

Cache-Oblivious Algorithms

• Cache-Oblivious Sorting for Burrows-Wheeler Transform (Advait D.Karande and Sriram Saroop)

• New Cache-Oblivious Algorithms (Zhang Jiahui and Neel Kamal)

• Cache-Oblivious, Lock-Free Algorithms (Seth Gilbert)